# Convolutional Neural Network NLP Natural Language Processing

### v1

@@ 2.2.2 Average Pooling
$t_{\texttt{avg}}\left(\mathbf{I},c,h,w\right)=\frac{1}{|\mathcal{N}_{K}(h,w)|}\sum_{(i,j)\in\mathcal{N}_{k}(h,w)}\mathbf{I}(c,i,j)$ (3)
@@ 2.2.3 Convolution
Convolution provides with its locality a favourable inductive bias for data on regular grids like images with highly correlated features in neighbourhoods. By optimizing its kernel parameters $\mathbf{K}\in\mathbb{R}^{C\times C\times K\times K}$ by gradient descent [39], it has been responsible for the breakthrough of Deep Learning in computer vision, providing a higher efficiency through its sparsity and reuse of parameter when compared to previously standard fully connected models. All this makes convolution a natural choice as a token mixer in image-based tasks.
$t_{\texttt{conv}}\left(\mathbf{I},\mathbf{K},c_{\texttt{out}},h,w\right)=\sum_{c=1}^{C}\sum_{(i,j)\in\mathcal{N}_{K}(h,w)}\mathbf{I}(c,i,j)\mathbf{K}(c_{\texttt{out}},c,i,j)$ (4)
Due to its connectivity over the channels, convolution has a high computational complexity of $\mathcal{O}(NK^{2}C^{2})$ and uses $K^{2}C^{2}$ parameters, mostly limiting the choice of kernel size $K=3$ for all but the first layers [37, 38].
@@ 2.2.4 Grouped Convolution
To overcome the high complexity of convolution, grouped convolution can be used [37]. Using $G$ filter groups, an output channel no longer depends on all input channels but on a subset of $\frac{C}{G}$. Thus, the depth connectivity can be reduced ($G<C$) or omitted ($G=C$). Note that $G$ has to be chosen so that $C$ is an integer multiple of $G$. To show the extrema, we chose $G=C$, shrinking the kernel weight matrix to $\mathbf{K}\in\mathbb{R}^{C\times K\times K}$ and thus omit the sum over the channel (compare Eq. 4)
$t_{\texttt{grouped\_conv}}\left(\mathbf{I},\mathbf{K},c,h,w\right)=\sum_{(i,j)\in\mathcal{N}_{K}(h,w)}\mathbf{I}(c,i,j)\mathbf{K}(c,i,j)$ (5)
This reduces complexity to $\mathcal{O}(NK^{2}C)$ and $K^{2}C$ parameters. Since channels are processed independently, grouped convolution with $G=C$ lacks channel cross-talk. However, as each MetaFormer block includes a channel-MLP (Fig. 1(b)) introducing cross-talk, we argue this limitation is negligible. This design parallels the depthwise separable convolution of MobileNet [40].
@@ 2.2.5 Self-Attention
With Self-Attention, Transformers have set a new state-of-the-art in natural language processing [6] and have also been well adapted to image processing tasks [7, 8, 9]. However, [23] provides evidence that the self-attention mechanism may not be the most important part of a successful model architecture for computer vision. To test this observation in the medical imaging domain, we include self-attention as a baseline for token mixing.
Let $\mathbf{p}_{r}=\mathbf{I}(h,w)\in\mathbb{R}^{C}$ define the current pixel’s feature vector with $r\leq N$ as the ravelled, flat index of $(h,w)$. For self-attention as token mixer $t_{\texttt{global\_attn}}$ each pixel $\mathbf{p}_{r}\in\mathbb{R}^{C}$ is transformed into the three roles: key $\mathbf{k}_{r}=\mathbf{W}_{k}\mathbf{p}_{r}$, value $\mathbf{v}_{r}=\mathbf{W}_{v}\mathbf{p}_{r}$, and query $\mathbf{q}_{r}=\mathbf{W}_{q}\mathbf{p}_{r}$, where each $\mathbf{W}\in\mathbb{R}^{C\times C}$ describes a learnable weight matrix. The transformed value of $\mathbf{p}_{r}$ is obtained by
$\mathbf{z}_{m,r}$ $=\sum_{s=1}^{N}\underset{s=1,\dots,N}{\texttt{softmax}}\left(\frac{\mathbf{q}_{m,r}^{\intercal}\mathbf{k}_{m,s}}{\sqrt{D}}\right)\cdot\mathbf{v}_{m,s},\ m=1,\dots,M$ (6)
@@
where $\mathbf{k},\mathbf{v}$, and $\mathbf{q}$ is split up into $M$ vectors of size $D$, forming attention heads $m$, allowing to learn different attention patterns. Their results are finally unified to $\mathbf{z}_{r}$ by $\mathbf{W}_{u}\in\mathbb{R}^{C\times C}$. In our experiments, we adapt the number of heads $M=\frac{C}{16}$ dynamically for each stage, dividing the number of channel $C$ by 16. Since self-attention is permutation-invariant, but images have an informative locality structure, we add positional embedding following [23]. While self-attention’s parameter count $4C^{2}$ is independent of the amount of processed pixels, its computational complexity quadratically grows with their amount $\mathcal{O}(N^{2}C)$. This significantly limits the applicability of self-attention for large input resolutions.
@@ 2.2.6 Local Self-Attention
To break up the quadratic runtime complexity and to infuse self-attention with the same advantages of convolution’s locality (e.g. inductive bias and translation equivariance), we limit the number of keys a query can attend to its direct neighbourhood. For that, we adapt the neighbourhood function’s signature $\mathcal{N}_{K}(r)$ to work with the flat index $r$ (cf. Eq. (1)). We extend Eq. (6) to construct $t_{\texttt{local\_attn}}$ as follows:
$\mathbf{z}_{m,r}=\sum_{s\in\mathcal{N}_{K}(r)}\underset{s\in\mathcal{N}_{K}(r)}{\texttt{softmax}}\left(\frac{\mathbf{q}_{m,r}^{\intercal}\mathbf{k}_{m,s}}{\sqrt{D}}\right)\cdot\mathbf{v}_{m,s},\ m=1,\dots,M.$ (8)
We implement $\mathcal{N}_{k}(r)$’s attention pattern using PyTorch’s flexattention framework 111pytorch.org/blog/flexattention. By restricting the receptive field of $\mathbf{q}$, local self-attention reduces complexity to $\mathcal{O}(N\cdot K^{2}\cdot C^{2})$ for typical resolutions ($K^{2}\ll N=HW$) compared to global self-attention, while keeping the parameter count unchanged at $(4\cdot C^{2})$.
Table 1: Runtime complexity in $\mathcal{O}$ notation, Floating Point Operations (FLOPs), and trainable parameter count of different token mixers in the MetaFormer block (including channel-MLP) regarding the kernel size $K$, number of in- and output channels $C$, and spatial input size $N=H\cdot W$. Note that in MetaFormer, no token mixer changes the spatial resolution or number of feature channels. Similarly, we only consider kernels of shape $(K\times K)$ comprising $K^{2}$ elements. FLOPs formulas are constructed ignoring technical optimisations in hard- and software.
@@ 3 Experiments
This section covers our experimental setup, describing the datasets and training routines. Since medical image datasets are often smaller than natural image datasets like ImageNet, we employ the smallest MetaFormer variant S12 with about 11.4 million trainable parameters (excluding the token mixers) to carry out our experiments. It comes with 64, 128, 320 and 512 channels for its four different stages. Each stage comprises two MetaFormer blocks, except for the third stage (six blocks, see Fig. 1(c)). For both settings, classification and segmentation, we compare the different token mixers introduced in Sec. 2.2 utilizing different pooling or kernel sizes where applicable (i.e. not for global attention and identity). Our code is available on GitHub222github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
@@ 3.1 Classification
For classification, we extend the MetaFormer encoder with a classification head, averaging the final feature map to a vector via Global Average Pooling (GAP) and then linearly mapping it to the number of classes (see Fig. 1(c)). For optimization, we employ AdamW [41] with a weight decay of $0.1$ using the default learning rate of $1e^{-3}$, which is reduced by a single cycle of cosine annealing [42] (timm implementation) to a minimum of $1e^{-5}$. We found it beneficial to employ a linear learning rate warm-up during the first 5 epochs of training. We adapt the number of epochs to dataset size, targeting 35k iterations with batch size 128. The final model is selected based on the highest validation F1-score measured at the end of each training epoch. For the loss function, we chose cross-entropy with label smoothing of 0.1 and class weighting by the inverse root class frequency (clamped to a maximum of 10). Data augmentation consists of random affine transformations with a transformation matrix sampled from a zero-mean normal distribution with 0.1 standard deviation. For additional regularisation, we follow [23] utilizing stochastic depth [43] and LayerScale [44]. All images were bilinearly reshaped to $224\times 224$.
Given the scarcity of medical data, pretrained weights are often essential. This applies especially for self-attention, which converges slower and requires more data due to its weaker inductive bias compared to convolution [45]. Consequently, we evaluate two MetaFormer variants: The first variant is trained from scratch, with the token mixer of all stages change to the specific experiment setting (denoted as $[T,T,T,T]$). Since there are no pretrained weights for all our deployed token mixers, we adapted the hybrid architecture of pooling, pooling, global self-attention, global self-attention, for which ImageNet weights have been made public [23], as our second pretrained variant. Here we keep the pooling in the first two stages and exchange the self-attention in the last two stages according to the experiment setting (denoted as $[P,P,T,T]$) and fine-tune the channel-MLP of all stages. Notably, the third stage contains six layers (vs. two in the others), so modifying it changes most of the token mixers. To measure the improvement provided by the pretrained weights, we also train the second variant ($[P,P,T,T]$) from scratch (result shown in appendix Tab. A.3).
@@ 3.1.1 Datasets
We include various classification datasets with vastly different numbers of training images from multiple medical image modalities using the MedMNIST V2 collection [46]. From the collection, we select tasks that are harder to classify and thus suitable for comparing token mixers. We gauge the difficulty of the twelve MedMNIST2D datasets by training a ResNet18 with the aforementioned routine (Sec. 3.1) for up to 500 epochs or 35k35\,k iterations, without a learning rate scheduler. We then select datasets with a validation F1-score below 0.95 for our studies. This procedure yielded four diverse datasets: PathMNIST [47] provides patches from hematoxylin and eosin stained colorectal cancer histological images for the task of classifying nine different types of tissues. The images are rotation invariant and are split into 89 996/10 004/7 18089\,996/10\,004/7\,180 train/validation/test images, making it the largest dataset in our work. DermaMNIST [48, 49] consists of dermatologic images of nine common pigmented skin lesions, leaving it also rotation invariant. The images are divided into 7 007/1 003/2 0057\,007/1\,003/2\,005 splits. With PneumoniaMNIST [50], we include pneumonia detection in chest X-rays with 4 708/524/6244\,708/524/624 images, respectively. Lastly, in OrganSMNIST [51, 52] we classify eleven organs in the sagittal view of 201 CT scans, obtaining 13 932/2 452/8 82713\,932/2\,452/8\,827 images respectively.
To measure the factor of a potential domain shift between the pretrained ImageNet weights and the medical data, we also include a small subset of the ImageNet dataset, since a full-size ImageNet training would have been unfeasible considering the number of experiments and available hardware. We settle on ImageWoof [53]: A non-trivial dataset of ten different dog breeds, coming with a predefined split of 9 035/3 9399\,035/3\,939 train/test images.
@@ 3.1.2 CNN Baseline
As a CNN baseline, we employ ResNet18 [38] (torchvision implementation), which matches the parameter count of MetaFormerS12. For models trained from scratch, we test multiple kernel sizes as in MetaFormer, while for pretrained models we use only kernel size 3, the only available pretrained variant.
@@ 3.2 Semantic Segmentation
For semantic segmentation, we extract the last block’s feature map of all four stages and employ the same light weighted all-MLP decoder used in the SegFormer [8] adding $525\,\mathrm{k}$ parameters to the MetaFormer. We chose this small decoder to minimise its impact on the segmentation tasks and have the MetaFormer be the main feature extractor, amplifying the potential performance difference between types of token mixers. Fig. 1(c) shows a schematic depiction of the decoder. The SegFormer decoder takes the feature maps from the four MetaFormer stages, equalizes the number of channels to 256 using linear layers, and upsamples them bilinearly to the first stage spatial resolution. A final two-layer channel-MLP with 256 hidden neurons then projects the concatenated $4\cdot 256$-channel feature map to the number of classes. Since, the feature map from the first stage is downsampled by factor four by the MetaFormer’s first patch embedding, the logits of the final predicted segmentation mask are bilinearly upsampled by factor four. It is worth noting, the MLP-decoder only includes pointwise operations and thus no token mixing takes place except within the bilinear upsampling (analogous to distance-weighted average pooling). We use the same training routine as for classification (see Sec. 3.1), only swapping the loss function to cross-entropy plus Dice loss (equally weighted) and setting a fixed number of epochs of 1000 regardless of the dataset size. Whenever a multiclass segmentation task includes the background class, we ignore it during the loss calculation.

### v2

@@ 3.2.2 Average Pooling
$t_{\texttt{avg}}\left(\mathbf{I},c,h,w\right)=\frac{1}{K^{2}}\sum_{i=-K}^{K}\sum_{j=-K}^{K}\mathbf{I}(c,h+i,w+j)$ (2)
@@ 3.2.3 Convolution
Convolution provides with its locality a favourable inductive bias for data on regular grids like images with highly correlated features in neighbourhoods. By optimizing its kernel parameters $\mathbf{K}\in\mathbb{R}^{C\times C\times K\times K}$ by gradient descent (LeCun et al., 1989), it has been responsible for the breakthrough of Deep Learning in computer vision, providing a higher efficiency through its sparsity and reuse of parameter when compared to previously standard fully connected models. All this makes convolution a natural choice as a token mixer in image-based tasks.
$t_{\texttt{conv}}$ $\left(\mathbf{I},\mathbf{K},c_{\texttt{out}},h,w\right)$ (3)
$=\sum_{c=1}^{C}\sum_{i=-K}^{K}\sum_{j=-K}^{K}\mathbf{I}(c,i,j)\mathbf{K}(c_{\texttt{out}},c,h+i,w+j)$
Due to its connectivity over the channels, convolution has a high computational complexity of $\mathcal{O}(NK^{2}C^{2})$ and uses $K^{2}C^{2}$ parameters, mostly limiting the choice of kernel size to $K=3$ for all but the first layers (Krizhevsky et al., 2012; He et al., 2016).
@@ 3.2.4 Grouped Convolution
To overcome the high complexity of convolution, grouped convolution can be used (Krizhevsky et al., 2012). Using $G$ filter groups, an output channel no longer depends on all input channels but on a subset of $\frac{C}{G}$. Thus, the depth connectivity can be reduced ($G<C$) or omitted ($G=C$). Note that $G$ has to be chosen so that $C$ is an integer multiple of $G$. To show the extrema, we chose $G=C$, shrinking the kernel weight matrix to $\mathbf{K}\in\mathbb{R}^{C\times K\times K}$ and thus omit the sum over the channel (compare Eq. 3)
$t_{\texttt{group\_conv}}$ $\left(\mathbf{I},\mathbf{K},c,h,w\right)$ (4)
$=\sum_{i=-K}^{K}\sum_{j=-K}^{K}\mathbf{I}(c,i,j)\mathbf{K}(c,h+i,w+j)$
This reduces complexity to $\mathcal{O}(NK^{2}C)$ and $K^{2}C$ parameters. Since channels are processed independently, grouped convolution with $G=C$ lacks channel cross-talk. However, as each MetaFormer block includes a channel-MLP (Fig. 1) introducing cross-talk, we argue this limitation is negligible. This design parallels the depthwise separable convolution of MobileNet (Howard et al., 2017).
@@ 3.2.5 Self-Attention
With Self-Attention, Transformers have set a new state-of-the-art in natural language processing (Vaswani et al., 2017) and have also been well adapted to image processing tasks (Dosovitskiy et al., 2021; Xie et al., 2021; Carion et al., 2020). However, Yu et al. (2022) provide evidence that the self-attention mechanism may not be the most important part of a successful model architecture for computer vision. To test this observation in the medical imaging domain, we include self-attention as a baseline for token mixing.
Let $\mathbf{p}_{r}=\mathbf{I}(h,w)=\mathbf{I}(\lfloor\tfrac{r}{W}\rfloor,r\mod W)\in\mathbb{R}^{C}$ define the current pixel’s feature vector with $r=h\cdot W+w\leq N$ as the ravelled, flat index of $(h,w)$. For self-attention as token mixer $t_{\texttt{global\_attn}}$, each pixel $\mathbf{p}_{r}\in\mathbb{R}^{C}$ is transformed into the three roles: key $\mathbf{k}_{r}=\mathbf{W}_{k}\mathbf{p}_{r}$, value $\mathbf{v}_{r}=\mathbf{W}_{v}\mathbf{p}_{r}$, and query $\mathbf{q}_{r}=\mathbf{W}_{q}\mathbf{p}_{r}$, where each $\mathbf{W}\in\mathbb{R}^{C\times C}$ describes a learnable weight matrix. The transformed value of $\mathbf{p}_{r}$ is obtained by
$\mathbf{z}_{m,r}$ $=\sum_{s=1}^{N}\underset{s=1,\dots,N}{\texttt{softmax}}\left(\frac{\mathbf{q}_{m,r}^{\intercal}\mathbf{k}_{m,s}}{\sqrt{D}}\right)\cdot\mathbf{v}_{m,s}$ (5)
@@
where $\mathbf{k},\mathbf{v}$, and $\mathbf{q}$ is split up into $M$ vectors of size $D$, forming attention heads $m=1,\dots,M$, allowing to learn different attention patterns. Their results are finally unified to $\mathbf{z}_{r}$ by $\mathbf{W}_{u}\in\mathbb{R}^{C\times C}$. In our experiments, we adapt the number of heads $M=\frac{C}{16}$ dynamically for each stage, dividing the number of channels $C$ by 16. Since self-attention is permutation-invariant, but images have an informative locality structure, we add positional embedding following Yu et al. (2022). While self-attention’s parameter count $4C^{2}$ is independent of the amount of processed pixels, its computational complexity quadratically grows with their amount $\mathcal{O}(N^{2}C)$. This significantly limits the applicability of self-attention for large input resolutions.
@@ 3.2.6 Local Self-Attention
To break up the quadratic runtime complexity and to infuse self-attention with the same advantages of convolution’s locality (e.g. inductive bias and translation equivariance), we limit the number of keys a query can attend to its direct neighbourhood. Let $\mathcal{N}$ define a function that returns the neighbouring pixel coordinates of a given location:
$\mathcal{N}_{K}(h,w)=\left\{\begin{aligned} i,j\in\mathbb{N}_{0}:\;&(|i-h|<\tfrac{K}{2})\\ &\land(|j-w|<\tfrac{K}{2})\end{aligned}\right\},$ (7)
where $K$ describes the kernel size in height and width, since we limit ourselves to quadratic kernels. We equivalently denote the neighbourhood function as $\mathcal{N}_{K}(r)$ with the flat index $r$. With it, we extend Eq. (6) to construct $t_{\texttt{local\_attn}}$ as follows:
$\mathbf{z}_{m,r}=\sum_{s\in\mathcal{N}_{K}(r)}\underset{s\in\mathcal{N}_{K}(r)}{\texttt{softmax}}\left(\frac{\mathbf{q}_{m,r}^{\intercal}\mathbf{k}_{m,s}}{\sqrt{D}}\right)\cdot\mathbf{v}_{m,s}.$ (8)
We implement $\mathcal{N}_{K}(r)$’s attention pattern using PyTorch’s flexattention framework111pytorch.org/blog/flexattention. By restricting the receptive field of $\mathbf{q}$, local self-attention reduces complexity to $\mathcal{O}(NK^{2}C)$ for typical resolutions ($K^{2}\ll N=HW$) compared to global self-attention (cf. Fig. 2(a)), while keeping the parameter count unchanged at $4C^{2}$.
Table 1: Runtime complexity in $\mathcal{O}$ notation, Floating Point Operations (FLOPs), and trainable parameter count of different token mixers in the MetaFormer block (plus channel-MLP) regarding the kernel size $K$, number of in- and output channels $C$, and spatial input size $N=H\cdot W$. Note that in MetaFormer, no token mixer changes the spatial resolution or number of feature channels. Similarly, we only consider kernels of shape $(K\times K)$ comprising $K^{2}$ elements. FLOPs formulas are constructed ignoring technical optimisations in hard- and software.
@@ 4 Experiments
This section covers our experimental setup, describing the datasets and training routines. Since medical image datasets are often smaller than natural image datasets like ImageNet, we employ the smallest MetaFormer variant S12 with about 11.4 million trainable parameters (excluding the token mixers) to carry out our experiments. It comes with 64, 128, 320 and 512 channels for its four different stages. Each stage comprises two MetaFormer blocks, except for the third stage (six blocks, see Fig. 2(b)). For both settings, classification and segmentation, we compare the different token mixers introduced in Sec. 3.2 utilizing different pooling or kernel sizes where applicable (i.e. not for global attention and identity). Our code is available on GitHub222github.com/multimodallearning/MetaFormerMedImaging/tree/clean_code.
@@ 4.1 Classification
For classification, we extend the MetaFormer encoder with a classification head, averaging the final feature map to a vector via Global Average Pooling (GAP) and then linearly mapping it to the number of classes (see Fig. 2(b)). For optimization, we employ AdamW (Loshchilov and Hutter, 2019) with a weight decay of $0.1$ using the default learning rate of $1e^{-3}$, which is reduced by a single cycle of cosine annealing (Loshchilov and Hutter, 2017) (timm implementation) to a minimum of $1e^{-5}$. We found it beneficial to employ a linear learning rate warm-up during the first 5 epochs of training. We adapt the number of epochs to dataset size, targeting 35k iterations with batch size 128. The final model is selected based on the highest validation F1-score measured at the end of each training epoch. For the loss function, we chose cross-entropy with label smoothing of 0.1 and class weighting by the inverse root class frequency (clamped to a maximum of 10). Data augmentation consists of random affine transformations with a transformation matrix sampled from a zero-mean normal distribution with 0.1 standard deviation. For additional regularisation, we follow (Yu et al., 2022) utilizing stochastic depth (Huang et al., 2016) and LayerScale (Touvron et al., 2021). All images were bilinearly reshaped to $224\times 224$.
Given the scarcity of medical data, pretrained weights are often essential. This applies especially for self-attention, which converges slower and requires more data due to its weaker inductive bias compared to convolution (Lu et al., 2022). Consequently, we evaluate two MetaFormer variants: The first variant is trained from scratch, with the token mixer of all stages change to the specific experiment setting (denoted as $[T,T,T,T]$). Since there are no pretrained weights for all our deployed token mixers, we adapted the hybrid architecture of pooling, pooling, global self-attention, global self-attention, for which ImageNet weights have been made public (Yu et al., 2022), as our second pretrained variant. Here we keep the pooling in the first two stages and exchange the self-attention in the last two stages according to the experiment setting (denoted as $[P,P,T,T]$) and fine-tune the channel-MLP of all stages. Notably, the third stage contains six layers (vs. two in the others), so modifying it changes most of the token mixers. To measure the improvement provided by the pretrained weights, we also train the second variant ($[P,P,T,T]$) from scratch (result shown in appendix Tab. 6).
@@ 4.1.1 Datasets
We include various classification datasets with vastly different numbers of training images from multiple medical image modalities using the MedMNIST V2 collection Yang et al. (2023). From the collection, we select tasks that are harder to classify and thus suitable for comparing token mixers. We gauge the difficulty of the twelve MedMNIST2D datasets by training a ResNet18 with the aforementioned routine (Sec. 4.1) for up to 500 epochs or 35k35\,k iterations, without a learning rate scheduler. We then select datasets with a validation F1-score below 0.95 for our studies. This procedure yielded four diverse datasets: PathMNIST (Kather et al., 2019) provides patches from hematoxylin and eosin stained colorectal cancer histological images for the task of classifying nine different types of tissues. The images are rotation invariant and are split into 89 996/10 004/7 18089\,996/10\,004/7\,180 train/validation/test images, making it the largest dataset in our work. DermaMNIST (Tschandl et al., 2018; Codella et al., 2019) consists of dermatologic images of nine common pigmented skin lesions, leaving it also rotation invariant. The images are divided into 7 007/1 003/2 0057\,007/1\,003/2\,005 splits. With PneumoniaMNIST (Kermany et al., 2018), we include pneumonia detection in chest X-rays with 4 708/524/6244\,708/524/624 images, respectively. Lastly, in OrganSMNIST (Bilic et al., 2019; Xu et al., 2019) we classify eleven organs in the sagittal view of 201 CT scans, obtaining 13 932/2 452/8 82713\,932/2\,452/8\,827 images respectively.
To measure the factor of a potential domain shift between the pretrained ImageNet weights and the medical data, we also include a small subset of the ImageNet dataset, since a full-size ImageNet training would have been unfeasible considering the number of experiments and available hardware. We settle on ImageWoof (Howard, 2019): A non-trivial dataset of ten different dog breeds, coming with a predefined split of 9 035/3 9399\,035/3\,939 train/test images.
@@ 4.1.2 CNN Baseline
As a CNN baseline, we employ ResNet18 (He et al., 2016) (torchvision implementation), which matches the parameter count of MetaFormerS12. For models trained from scratch, we test multiple kernel sizes as in MetaFormer, while for pretrained models we use only kernel size 3, the only available pretrained variant.
@@ 4.2 Semantic Segmentation
For semantic segmentation, we extract the last block’s feature map of all four stages and employ the same light weighted all-MLP decoder used in the SegFormer (Xie et al., 2021) adding $525\,\mathrm{k}$ parameters to the MetaFormer. We chose this small decoder to minimise its impact on the segmentation tasks and have the MetaFormer be the main feature extractor, amplifying the potential performance difference between types of token mixers. Fig. 2(b) shows a schematic depiction of the decoder. The SegFormer decoder takes the feature maps from the four MetaFormer stages, equalizes the number of channels to 256 using linear layers, and upsamples them bilinearly to the first stage spatial resolution. A final two-layer channel-MLP with 256 hidden neurons then projects the concatenated $4\cdot 256$-channel feature map to the number of classes. Since, the feature map from the first stage is downsampled by factor four by the MetaFormer’s first patch embedding, the logits of the final predicted segmentation mask are bilinearly upsampled by factor four. It is worth noting, the MLP-decoder only includes pointwise operations and thus no token mixing takes place except within the bilinear upsampling (analogous to distance-weighted average pooling). We use the same training routine as for classification (see Sec. 4.1), only swapping the loss function to cross-entropy plus Dice loss (equally weighted) and setting a fixed number of epochs of 1000 regardless of the dataset size. Whenever a multiclass segmentation task includes the background class, we ignore it during the loss calculation.